OpenStudy (anonymous):

y = 3 sin(2x) In this, would the 2 be the period?

OpenStudy (psymon):

2x = 2(pi) (normal sin period) So that means the period of 2x would be solving for x. So divide both sides by 2 and you get simply pi.

OpenStudy (anonymous):

...What?

OpenStudy (psymon):

Sorry. Im sure sami has a better answer for ya.

OpenStudy (anonymous):

period of sin(x) is 2pi generally if \[\Large \sin(ax)\] then period will be \[\Large \frac{2\pi}{a}\] here in 3sin(2x) a=2 can you find it now ?

OpenStudy (anonymous):

...No. I'm even more confused.

OpenStudy (jdoe0001):

hehhe

OpenStudy (psymon):



OpenStudy (anonymous):

let me give you few examples just remember the formula if \[\Large \sin(ax)\] period is given by \[\Large \frac{2\pi}{a}\] Examples 1) period of sin(5x) is \[\Large \frac{2\pi}{5}\] because a=5 here 2) period of sin(3x) \[\Large \frac{2\pi}{3}\] because a=3 here 3) period of sin(7x) is \[\Large \frac{2\pi}{7}\] because a=7 here can you now find the period of 3sin(2x) ?

OpenStudy (jdoe0001):



OpenStudy (anonymous):

(2pi)/2???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or you can say pi as 2 will be cancellled with 2

OpenStudy (anonymous):

Ohhh. What about amplitudes?

OpenStudy (anonymous):

amplitute is affeted by the constant multiplied with it for example bsin(ax) here period is 2pi/a and amplitude is b so 1) 2sin(5) amplitude is 2 2) 3sin(7x) amplitude is 3 3) 6 sin(2x) amplitude is 6 so just remember whats inside the parenthesis contributes to period and the constant bieng multiplied decide the amplitude can you now find the amplitude of 3sin(2x) ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

What if you wanna find a period of something that is like y = 1/3cos(x/2 )?